Assume there are three companies that in the past year paid exactly the same annual dividend of S$1 52 a share In addition, the future annual rate of growth in dvidends for each of the three companies has been estimated as follows: EEB Assume also that as the result of a strange set of circumstances, these three companies all have the same required rate of retum (r: 11%) a. Use the appropriate DVM to value each of these companies b. Comment briefly on the comparative values of these three companies What is the major cause of the differences among these three valuations? a. For Buggies-Are-Us, the value of the companys common shares is (Round to the nearest cent) For Steady Freddie, Inc, the value of the companys common shares is s(Round to the nearest cent) For Gang Buster Group, the value of the companys common shares is s ll (Round to the nearest cent) b. Comment briefly on the comparative values of these three companies, What is the major cause of the differences among these three valuations? (Select frhe best choice below) O A. The value of Buggies Are-Us is $13 82 compared to S40 75 for Steady Freddie, Inc caused by the difference in dividend growth rates. The Buggles-Are-Us dividends do not grow, resulting in the lowest value. The dividends of Steady Freddie, inc grow at a constant rate of 7% forever, whereas Gang Busters Groups di dends grow at approxmately 12% for the first four years and 7% from year five to the foreseeable future The higher growth in dividends in the earlier years causes the stock of Gang Busters Group to be worth more than the Steady Freddie, Inc, stock The value of Buggies Are Us is $13 82, compared to S40 75 for Steady Freddie Inc caused by the difference in dividend growth rates The Buggles-Are-Us dividends do not grow, resulting in the lowest value. The dividends of Steady Freddie Inc. grow at a constant rate of 7% forever, whereas Gang Busters Groups dividends gro at approximately 12% for the first four years an and S49 10 for Gang Busters Group The difference in values is O B. and S49 10 for Gang Busters Group The difference in values is
here are three companies that in the past year paid exactly the same annual dividend of $1.52 a share. In addition, the future annual rate of growth in for each of the three companies has been estimated as follows: EEB Assume also that as the result of a strange set of circumstances, these three s all have the same required rate of return (r= 11%) e appropriate DVM to value ent brielly on the comparat ggies-Are-Us, the value of y Freddie, Inc, the value o Buster Group, the value of briefly on the comparat -X ree valuations? Data Table (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a spreadsheet) Buggies-Are-Us Steady Freddie, Inc g-7% for the foreseeable future) Gang Buster Group Year 1 Year 2 Year 3 Year 4 ent briefly on the compara low) ree valuations? (Selečt the b (ie, dividends are expected to remain at $1.52/share) $1.92 $2.16 $2.43 796 he value of Buggies-Are-Us used by the difference in d eddie, Inc, grow at a const om year five to the foreseea e Steady Freddie, Inc, stod he value of Buggies-Are-Us used by the difference in dl reddie, Inc, grow at a constar he difference in values is The dividends of Steady or the first four years and 7% s Group to be worth more than Year 5 and beyond Print Done he difference in values is The dividends of Steady for the first four years and 11%
Answer #1

Price of the stock = D(1+g) / (r-g)

where D= Dividend, g= growth rate , r = rate of return

a. Buggies are us stock price = 1.52 (1+0) /(11%-0) = $13.82

Steady Freddie Inc. = 1.52(1+7%) /(11%-7%) = $40.66

Gang Buster Group is calculated as below:

Year Dividend PVIF11% Present Value
[Dividend * PVIF11%]
1 1.71 0.9091 1.55
2 1.92 0.8264 1.59
3 2.16 0.7513 1.62
4 2.43 0.6830 1.66
Total 6.42

Value of the stock at year 5 = 2.43(1+7%) (11%-7%) = $65.0025

Present value of the stock at the end of 5th year = 65.0025* 0.683 = $44.39

Value of the stock = $44.39+ $6.42 = $50.82 ( rounded off)

b) The correct option is (A).